# Quasi-aperiodic grain boundary phases of tilt grain boundaries in refractory metals

### v2

@@ Abstract
We report new quasi-aperiodic, ground-state structures and phase transitions in 5[001] tilt grain boundaries (GBs) in body-centered cubic (BCC) refractory metals Nb, Ta, Mo, and W. 5 tilt GBs have been extensively investigated over the past several decades, with their ground-state structure—composed of kite-shaped structural units—previously thought to be well understood. By performing a rigorous GB structure search that optimizes the number of atoms in the boundary core, we predict new quasi-aperiodic “split kite” phases analogous to those previously found in GBs in face-centered cubic metals. Our results suggest that complex aperiodic phases of GBs appear to be a general phenomenon. Density functional theory calculations yield similar GB energies and structures for kite and split kite phases in all BCC metals studied. Phase-contrast image simulations of split kites show better agreement with experimental observations, offering an alternative explanation for previous microscopy results and motivating future atomically resolved imaging of the GB structure.
@@
More recently, improved simulation methodologies that allowed the number of atoms in the GB to change revealed that most empirical embedded-atom method (EAM) potentials of FCC metals predict a completely different ground-state structure, which was named “split kites” [20]. This structure required inserting or removing about half of the plane of atoms in the boundary core. Moreover, split kites exist in multiple energy-degenerate states characterized by different arrangements of their structural units without a well-defined periodic unit cell [21]. Due to its large area, the split kite structure in FCC metals has not been investigated with DFT calculations; however, subsequent studies of [110] symmetric tilt GBs in BCC metals revealed analogous complex aperiodic states that were confirmed by DFT calculations [22] and subsequently validated by direct observations in BCC Fe [23].
Motivated by these studies, we revisit tilt GBs in BCC refractory metals represented by Nb, Ta, Mo, and W. We perform GB structure search using a recently developed Grand canonical Interface Predictor (GRIP) code that optimizes the GB core configuration [24]. We find new aperiodic structures of these tilt boundaries that are analogous to the split kites observed in FCC metals. These complex structures are confirmed using DFT to be the ground states in Ta, Mo, and W, and are at least metastable in the other BCC systems studied.
Results: In this work, we focus on two symmetric tilt (310)[001] and (210)[001] boundaries with misorientation angles around the common tilt axis of and , respectively. The energy of the system is modeled using EAM and modified embedded-atom method (MEAM) interatomic potentials, as well as DFT calculations. Using GRIP, we systematically explore GB structures by uniformly sampling relative translations between the two grains, randomly removing a fraction of atoms from the boundary plane, and dynamically sampling different configurations using empirical potentials [24]. As described in our previous studies [20], we track the GB atomic density, , by calculating
@@
The generated BCC split kite phases appear to be aperiodic. Using GRIP, we explore larger area reconstructions of these GBs by tiling the smallest unit cell up to 25 times in the direction along the [001] tilt axis. The nearly degenerate lowest-energy structures for different reconstruction sizes of 1 6, 1 9 and 1 25 are shown in Figure 2a–c. While the structures viewed down the tilt axis appear nearly indistinguishable, the different reconstructions have distinct atomic arrangements along the tilt axis. When viewed from the side, the atoms belonging to the tips of the kites shown in yellow form a periodic wave with a period commensurate with the simulation cell size. Figure 2d shows the GB energy of the different split kite structures as a function of repetition along the tilt axis. The energy converges only for very large reconstructions of 10 and larger, although a structure with a similar low energy can be found for smaller reconstructions (e.g., 1 6 for Ta (210)), which we select for subsequent DFT calculations. Even for very large sizes, the energy plot shows oscillations in . A similar convergence behavior is observed in the values of for the minimum-energy structure at each reconstruction size in Figure 2d.
Split kite phases of 5(210) were identified by empirical potentials in all BCC metals we studied. However, for Mo, W, and Fe, their energies predicted by these models are 5– higher than those of the kite ground-state structure. Representative results are shown in Figure 3 for the Mo (210) tilt GB, and similar results for W and Fe are in Supp. Figure S1 and Supp. Figure S2, respectively [25]. Split kites are a metastable phase of this boundary; they are local minima with respect to and remain stable at finite temperatures (homologous temperature of ) in MD simulations with periodic boundary conditions.
To validate the stability of the split kite structures predicted at by GRIP and demonstrate possible GB phase transitions, we perform finite-temperature MD simulations at temperatures ranging from to 0.7. To model GB phase transitions from kites to split kites, the GBs were terminated at open surfaces following the methodology proposed in Ref. [20]. Open surfaces act as sources and sinks of atoms and effectively introduce a grand-canonical environment in the GB core. The simulations reveal that the split kite phases of all metals are dynamically stable. First-order phase transitions are observed with open-surface boundary conditions as illustrated in Figure 1g.
The use of empirical potentials (see Supp. Table S1) enables more thorough structure exploration using GRIP, including the searches of large-area reconstructions that are currently beyond the reach of DFT calculations. Previous studies have demonstrated the accuracy of empirical potentials in predicting new low-energy structures which were subsequently confirmed by DFT and observed in experiments [22, 23]. However, it is also well known that predicted GB structures and energies are sensitive to the choice of empirical potential [26]. We compare the different predictions of EAM and MEAM potentials in Supp. Figure S3 and Figure S4. With this in mind, we calculate the GB energies of each phase obtained from GRIP using DFT by relaxing a smaller block of atoms extracted from the GB region. Additional details on the methodology are provided in the Supplemental Material [25].
@@
Discussion: In this work, we revisit structures of [001] symmetric tilt GBs in refractory BCC metals and study their phase behavior using the recently developed GRIP tool [24]. Contrary to established and well-known structures, we find new ground states and demonstrate multiple GB phases by optimizing the number of atoms in the boundary core and exploring large-area reconstructions. The new BCC split kite phases in [001] tilt GBs can be realized in unit cells of different sizes and appear to be aperiodic (Figure 2), similar to previously reported split kite structures in [001] tilt GBs in FCC metals [20] and other [110] tilt GBs in BCC metals [22]. Our findings provide further evidence that complex aperiodic phases of GBs could be a general phenomenon.
@@
The BCC metals investigated here have different relative stability between the two GB phases and different ground-state structures, highlighting the potential importance of electronic effects. For the reconstructions accessible to DFT, we find that regular kites and split kites have close GB energies, and it is possible that larger area reconstructions could further lower the DFT energies. The existence of metastable split kite phases suggests that under some conditions they may be stabilized by temperature, pressure, or chemical segregation to become the ground states [27].
@@
Structures of GBs in several refractory metals were previously systematically investigated by high-resolution TEM (HRTEM) techniques [8, 9, 11, 10, 12, 13]. Unfortunately, in these earlier studies the GB structure was not observed directly. Instead, it was inferred by matching the simulated intensity patterns derived from structures predicted by atomistic modeling with the observed phase contrast in the HRTEM images. To compare our predicted structures with these previous experimental results, we generate the simulated images for split kite and regular kite structures using the multislice method implemented in abTEM [28]. The simulated images of the two different structures can be clearly distinguished in Figure 4. The wave pattern at the tip of the split kite structure introduces a characteristic broadening in the intensity that is notably absent in the regular kite. The same feature can be seen in the experimental image, suggesting that the split kite structure for Nb (310) may be a better match with previous HRTEM results and that our predicted structures may have already been experimentally observed.
@@
Previous experimental studies have also reported rigid body translations of (002) across the GB plane along the [001] tilt axis [29]. We report similar displacements for kites in Nb, Ta, and Mo in Supp. Table S3, consistent with some [16, 17] but not all [18] ab initio studies. Notably, the lack of translations in our simulated Nb (310) split kite structure, in agreement with experimental observations [9], further supports the possibility that this GB phase was previously observed. Given the sensitivity of the simulated structures and images to the simulation parameters, our results motivate further experimental studies using modern aberration-corrected electron microscopes that can directly resolve atomic positions.
@@ 2 Materials Science Division, Lawrence Livermore National Laboratory, Livermore, CA 94550, USA
To study GB phase stability and transitions, we perform finite-temperature MD simulations using methods adapted from previous work [24]. Briefly, we replicate the optimized GB structures in the - and -directions until the simulation cell is around in the -direction and in the -direction along the tilt axis. We freeze the bottom layer of atoms and constrain the top layer to be semi-rigid throughout the simulation, which lasts at least . We use periodic boundary conditions (PBCs) in the -direction and both PBCs and open surfaces with of vacuum in the -direction. We scan a range of temperatures between 0.2 and 0.7 of the melting temperature of each system. For clarity of visualization, we relax all structures at using a conjugate gradient minimization scheme.
Phase-contrast image simulations
To compare our results to previous experimental studies that used high-resolution transmission electron microscopy (HRTEM), we perform phase-contrast image simulations using the multislice algorithm implemented in the abTEM code [28]. The optimized grain boundary structures from GRIP for Nb (310) are used as input structures, replicated as necessary to achieve a lateral distance of at least and a thickness of at least along the tilt axis. We use 64 frozen phonons to approximate atomic vibrations [62] with a standard deviation of atomic displacements of and the infinite potential parametrization from Kirkland [63]. The electron beam energy is set at and the contrast transfer function parameters are (defocus of ), (spherical aberration), and .

### v3

@@ Abstract
We report new ground-state structures and phase transitions in 5[001] tilt grain boundaries (GBs) in body-centered cubic (BCC) refractory metals Nb, Ta, Mo, and W. 5 tilt GBs have been extensively investigated over the past several decades, with their ground-state structure—composed of kite-shaped structural units—previously thought to be well understood. By performing a rigorous GB structure search that optimizes the number of atoms in the boundary core, we predict new quasi-aperiodic “split kite” phases analogous to those previously found in GBs in face-centered cubic metals. Our results suggest that complex aperiodic phases of GBs appear to be a general phenomenon, as validated through density functional theory calculations. Moreover, the atoms in the split kite phase demonstrate distinct collective diffusion dynamics. Phase-contrast image simulations of split kites show better agreement with experimental observations, offering an alternative explanation for previous microscopy results and motivating future atomically resolved imaging of the GB structure.
@@
More recently, improved simulation methodologies that allowed the number of atoms in the GB to change revealed that most empirical embedded-atom method potentials of FCC metals predict a completely different ground-state structure, which was named “split kites” [20]. This structure required inserting or removing about half of the plane of atoms in the boundary core. Moreover, split kites exist in multiple energy-degenerate states characterized by different arrangements of their structural units without a well-defined periodic unit cell [21]. Due to its large area, the split kite structure in FCC metals has not been investigated with DFT calculations; however, subsequent studies of [110] tilt GBs in BCC metals revealed other complex aperiodic structures that were confirmed by DFT calculations [22] and subsequently validated by direct observations in BCC Fe [23]. Complex GB phases have demonstrated impacts on properties such as GB migration under shear [24] and GB diffusivity [25].
Motivated by these studies, we revisit tilt GBs in BCC refractory metals represented by Nb, Ta, Mo, and W. We perform GB structure search using a recently developed Grand canonical Interface Predictor (GRIP) code that optimizes the GB core configuration [26]. We discover new phases of these tilt boundaries that are analogous to the split kites observed in FCC metals. These complex structures are confirmed using DFT to be the ground states in Ta, Mo, and W, and are at least metastable in the other BCC systems studied.
Results: In this work, we focus on two symmetric tilt (310)[001] and (210)[001] boundaries with misorientation angles around the common tilt axis of and , respectively. The energy of the system is modeled using empirical potentials and DFT calculations. Using GRIP, we systematically explore GB structures by uniformly sampling relative translations between the two grains, randomly removing a fraction of atoms from the boundary plane, and dynamically sampling different configurations using empirical potentials [26]. As described in our previous studies [20], we track the GB atomic density, , by calculating
@@
The generated BCC split kite phases appear to be quasi-aperiodic. Using GRIP, we explore larger area reconstructions of these GBs by tiling the smallest unit cell up to 25 times in the direction along the [001] tilt axis. The nearly degenerate lowest-energy structures for different reconstruction sizes of 1 6, 1 9 and 1 25 are shown in Figure 2a–c. While the structures viewed down the tilt axis appear nearly indistinguishable, the different reconstructions have distinct atomic arrangements along the tilt axis. When viewed from the side, the atoms belonging to the tips of the kites shown in yellow form a periodic wave with a period commensurate with the simulation cell size. Figure 2d shows the GB energy of the different split kite structures as a function of repetition along the tilt axis. The energy converges only for very large reconstructions of 10 and larger, although a structure with a similar low energy can be found for smaller reconstructions (e.g., 1 6 for Ta (210)), which we select for subsequent DFT calculations. Even for very large sizes, the energy plot shows oscillations in . A similar convergence behavior is observed in the values of for the minimum-energy structure at each reconstruction size in Figure 2d. Additional top views of different cell sizes in the direction orthogonal to the tilt axis are shown in Supp. Figure S1. [27]
Split kite phases of 5(210) were identified by empirical potentials in all BCC metals we studied. However, for Mo, W, and Fe, their energies predicted by these models are 5– higher than those of the kite ground-state structure. Representative results are shown in Figure 3 for the Mo (210) tilt GB, and similar results for W and Fe are in Supp. Figure S2 and Supp. Figure S3, respectively. Split kites are a metastable phase of this boundary; they are local minima with respect to and remain stable at finite temperatures in MD simulations with periodic boundary conditions (PBCs).
To validate the stability of the split kite structures predicted at by GRIP and demonstrate possible GB phase transitions, we perform finite-temperature MD simulations at homologous temperatures ranging from to 0.7. To model GB phase transitions from kites to split kites, the GBs were terminated at open surfaces following the methodology proposed in Ref. [20]. Open surfaces act as sources and sinks of atoms and effectively introduce a grand-canonical environment in the GB core. First-order phase transitions are observed with open-surface boundary conditions as illustrated in Figure 1g and Supp. Movie 1.
Moreover, the simulations with PBCs reveal that the split kite phases of all metals are dynamically stable at high temperatures. These simulations also reveal that, despite the close energies of the two GB phases, the difference in their atomic structure has a significant impact on diffusivity and the diffusion mechanism. Figure 4 and Supp. Movie 2 show that the atoms forming the wave-like pattern in the split kites exhibit collective motion, a key component of short-circuit diffusion in GBs [28]. The wave is performing a random walk, moving the entire column of atoms and enabling large displacements of atoms along the tilt axis. These large displacements seem to be independent of the displacements in adjacent tip columns, as seen in Supp. Movie 3. In regular kites, on the other hand, the diffusion mechanism is represented by uncorrelated nearest-neighbor swaps and much slower diffusion (Supp. Movie 4).
@@
The use of empirical potentials (see Supp. Table S1) enables more thorough structure exploration using GRIP, including the searches of large-area reconstructions that are currently beyond the reach of DFT calculations. Previous studies have demonstrated the accuracy of empirical potentials in predicting new low-energy structures which were subsequently confirmed by DFT and observed in experiments [22, 23]. However, it is also well known that predicted GB structures and energies are sensitive to the choice of empirical potential [29]. We compare the different predictions of empirical potentials in Supp. Figure S4 and Figure S5. With this in mind, we calculate the GB energies of each phase obtained from GRIP using DFT by relaxing a smaller block of atoms extracted from the GB region. Additional details on the methodology are provided in the Supplemental Material [27].
@@
Discussion: In this work, we revisit structures of [001] symmetric tilt GBs in refractory BCC metals and study their phase behavior using the recently developed GRIP tool [26]. Contrary to established and well-known structures, we find new ground states and demonstrate multiple GB phases by optimizing the number of atoms in the boundary core and exploring large-area reconstructions. The new BCC split kite phases in [001] tilt GBs can be realized in unit cells of different sizes and periodicities (Figure 2), similar to previously reported split kite structures in [001] tilt GBs in FCC metals [20] and other [110] tilt GBs in BCC metals [22, 23]. In this context, our findings provide further evidence that complex aperiodic phases of GBs could be a general phenomenon.
@@
The BCC metals investigated here have different relative stability between the two GB phases and different ground-state structures. For the reconstructions accessible to DFT, we find that regular kites and split kites have close GB energies, and it is possible that larger area reconstructions could further lower the DFT energies. The existence of metastable split kite phases suggests that under some conditions they may be stabilized by temperature, pressure, or chemical segregation to become ground states [30].
@@
Structures of GBs in several refractory metals were previously systematically investigated by high-resolution TEM (HRTEM) techniques [8, 9, 11, 10, 12, 13]. Unfortunately, in these earlier studies the GB structure was not observed directly. Instead, it was inferred by matching the simulated intensity patterns derived from structures predicted by atomistic modeling with the observed phase contrast in the HRTEM images. To compare our predicted structures with these previous experimental results, we generate the simulated images for split kite and regular kite structures using the multislice method implemented in abTEM [31]. The simulated images of the two different structures can be clearly distinguished in Figure 5 (see Supp. Figure S7 for HAADF-STEM simulations). The wave pattern at the tip of the split kite structure introduces a characteristic broadening in the intensity that is notably absent in the regular kite. The same broadening can be seen in the experimental image, suggesting that our predicted split kite structure for Nb (310) may be a better match with previous HRTEM results and may have been experimentally observed.
@@
Previous experimental studies have also reported rigid body translations of (002) across the GB plane along the [001] tilt axis [32]. We report similar displacements for kites in Nb, Ta, and Mo in Supp. Table S3, consistent with some [16, 17] but not all [18] ab initio studies. Notably, the lack of translations in our simulated Nb (310) split kite structure is also in agreement with experimental reports [9]. Given the sensitivity of the simulated structures and images to the simulation parameters, our results motivate further experimental studies using modern aberration-corrected electron microscopes that can directly resolve atomic positions. Exact mechanisms leading to the formation of complex GB structures are currently unknown and warrant further investigation.
@@ 2 Materials Science Division, Lawrence Livermore National Laboratory, Livermore, CA 94550, USA
SUPPLEMENTARY MOVIE 1
Video of a first-order phase transformation in the Ta (210) GB at , where the regular kites transform into split kites at approximately 00:21 in the movie. The MD simulation is performed with open surfaces on both sides and each snapshot is relaxed for clarity of visualization. The movies can be found with the published manuscript or in Google Drive (relaxed, unrelaxed).
SUPPLEMENTARY MOVIE 2
Video of the side view (looking down ) of one atomic column of the tips of split kites. The quasi-aperiodic wave pattern persists at high temperature. A single atom is colored blue (tracer) to track the dynamics, which feature collective diffusion back and forth along the atomic column. Each frame is and the whole video is . The movie can be found with the published manuscript or in Google Drive.
SUPPLEMENTARY MOVIE 3
Video of the side view (looking down ) of two adjacent atomic columns of the tips of split kites. A single atom is colored in each atomic column (blue in the yellow wave and magenta in the green wave) to track the independent wave oscillations in separate GB cores. Each frame is and the whole video is . The movie can be found with the published manuscript or in Google Drive.
SUPPLEMENTARY MOVIE 4
Video of the side view (looking down ) of one atomic column of the tips of kites. A single atom is colored blue to track the dynamics, which are slow and feature uncorrelated nearest-neighbor hops. Each frame is and the whole video is . The movie can be found with the published manuscript or in Google Drive.
@@
To study GB phase stability and transitions, we perform finite-temperature MD simulations using methods adapted from previous work [26]. Briefly, we replicate the optimized GB structures in the - and -directions until the simulation cell is around in the -direction and in the -direction along the tilt axis. We freeze the bottom layer of atoms and constrain the top layer to be semi-rigid throughout the simulation, which lasts at least . We use periodic boundary conditions (PBCs) in the -direction and both PBCs and open surfaces with of vacuum in the -direction. We scan a range of temperatures between 0.2 and 0.7 of the melting temperature of each system. For clarity of visualization, we relax all structures at using a conjugate gradient minimization scheme. In Figure 4, the atomic positions are time-averaged across 1500 time steps () instead to preserve some high-temperature behavior.
Image simulations
To compare our results to previous experimental studies that used high-resolution transmission electron microscopy (HRTEM), we perform phase-contrast image simulations using the multislice algorithm implemented in the abTEM code [31]. The optimized grain boundary structures from GRIP for Nb (310) are used as input structures, replicated as necessary to achieve a lateral distance of at least and a thickness of at least along the tilt axis. We use 64 frozen phonons to approximate atomic vibrations [65] with a standard deviation of atomic displacements of and the infinite potential parametrization from Kirkland [66]. The electron beam energy is set at and the contrast transfer function parameters are (defocus of ), (spherical aberration), and . For high-angle annular dark-field STEM image simulations, the same electron beam energy was used with a semi-converge angle of , following Ref. [5].